a body projected in vertically upward direction is at same height after time t1 and t2 from start. the total time of flight is

Grade:

1 Answers

Prakhar Tayal
10 years ago

The answer is time=t1+t2.

How?

x=ut-.5gt2

2x=2ut-gt2

gt2 -2ut+2x=0

as t1 and t2 are roots of this equation hence 

t1+t2=2ut/g.   1 equation

Now consider a normal plot in where a ball comes back and we find its total flight time-

0=ut-1/2gt2

.5gt2=ut

.5gt=u

2u/g=t  2nd equation

from 1 and 2 nd equation

t=t1+t2

Regards,

Prakhar